Im having trouble trying to set up and solve this word problem.
The estimated number of insects on Earth is 1x10 to the 19th power. The number of people on Earth currently is estimated to be 6 x 10 to the 9th power. What is the estimate for the number of insects per person?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The estimated number of insects on Earth is 1x10 to the 19th power. The number of people on Earth currently is estimated to be 6 x 10 to the 9th power. What is the estimate for the number of insects per person?
---------------
insects per person (1x10^19)/(6x10^9) = (1/6)x(10^10) = 1,666,666,666
-------------
Cheers,
Stan H.
-------

Answer by Alan3354(69443)

To find bugs/person, divide the # of bugs by the # of people.